# Enhancing Urban Visual Place Recognition for Crowdsourced Flood Imagery via LLM-Guided Attention

### v1

@@ 3.2 Attention Integration into VPR models for Inference
The LLM-generated attention maps are integrated into pre-trained VPR models through a general principle: modulate spatial feature contributions in the aggregator according to LLM-derived importance maps while preserving the learned structure of the pre-trained model. This enables adaptive emphasis on semantically discriminative regions without retraining the backbone networks.
@@ 3.2.1 General Framework Formulation
Consider a pre-trained VPR model with backbone feature extractor and aggregator module $\mathcal{G}$. For a query image, the backbone produces a spatial feature map $\mathbf{X}\in\mathbb{R}^{C\times H\times W}$, where $C$ denotes the number of feature channels and $H\times W$ represents the spatial resolution. The standard inference pipeline computes the global descriptor as:
$\mathbf{v}=\mathcal{G}(\mathbf{X})\in\mathbb{R}^{D}$ (3)
where $\mathcal{G}$ aggregates spatial features into a $D$-dimensional global descriptor $\mathbf{v}$ for similarity matching against the reference database.
In the VPR-AttLLM framework, we augment this process by incorporating the LLM-generated attention map $\mathbf{A}^{\text{LLM}}\in\mathbb{R}^{H\times W}$ at the final spatial aggregation stage. To control the degree of LLM influence, we introduce a blending coefficient $\alpha\in[0,1]$ that interpolates between the original aggregation and full LLM-guided attention. The attention-modulated inference proceeds as:
@@
where $\mathbf{v}^{\text{LLM}}$ denotes the attention-weighted global descriptor, and $\alpha$ represents LLM impact ratio. When $\alpha=0$, the original pre-trained model behavior is recovered; when $\alpha=1$, full LLM-derived attention is applied. The attention map modulates spatial features within the aggregator $\mathcal{G}$ by scaling each spatial location’s contribution: regions emphasized by the LLM (where $A^{\text{LLM}}_{i,j}>1$) contribute more strongly to the final descriptor, while less discriminative areas (where $A^{\text{LLM}}_{i,j}<1$) are proportionally suppressed. All backbone parameters remain frozen—only the aggregation layer receives attention-modulated inputs during inference.
Critically, this attention integration applies only to query images at inference time. Reference database images are processed using the original pre-trained model without LLM augmentation. This asymmetric design offers two key advantages: (1) it adjusts the query descriptor formation to emphasize discriminative cues under challenging conditions (cross-source imagery, extreme weather, significant viewpoint variations), and (2) it avoids the computational overhead of processing the entire reference database through the LLM, which would be prohibitively expensive for large-scale deployment.
The specific integration mechanism adapts to the mathematical structure of each aggregator type while maintaining the general principle of attention-based spatial modulation. We demonstrate this across two predominant aggregation paradigms in state-of-the-art VPR architectures: GeM (Generalized Mean) pooling (Section 3.2.2) and cluster-based aggregation (Section 3.2.3). The overall VPR-AttLLM framework is illustrated in Figure 3.
3.2.2 LLM-guided Attention for GeM Pooling
GeM (Generalized Mean Pooling) [26] is widely used as the aggregation function before generating the final descriptor in many VPR models, such as CosPlace, EigenPlaces, CricaVPR. For a feature map $\mathbf{X}\in\mathbb{R}^{C\times H\times W}$, standard GeM pooling computes the output for each channel $c$ as:
$y_{c}=\left(\frac{1}{HW}\sum_{i=1}^{H}\sum_{j=1}^{W}x_{c,i,j}^{p}\right)^{1/p}$ (5)
where $p>1$ is a learnable parameter optimized during training that controls the trade-off between average and max pooling, and $x_{c,i,j}$ denotes the feature value at spatial position $(i,j)$ in channel $c$.
To understand how the pre-trained GeM aggregator implicitly weights different spatial regions, we infer the existing spatial contribution map from the trained parameter $p$. Specifically, for each spatial location $(i,j)$, the effective contribution to the pooled descriptor can be quantified as:
$w^{\text{GeM}}_{i,j}=\left(\frac{1}{HW}\sum_{c=1}^{C}x_{c,i,j}^{p}\right)^{1/p}$ (6)
normalized by its spatial mean to obtain a relative weight map $\mathbf{W}^{\text{GeM}}\in\mathbb{R}^{H\times W}$ that reflects the learned spatial emphasis of the original model.
We then compute a blended attention map that modulates the existing GeM contribution pattern with the LLM attention map $\mathbf{A}^{\text{LLM}}\in\mathbb{R}^{H\times W}$ using blending ratio $\alpha$. The final weight map is defined as:
$\mathbf{W}^{\text{final}}=\mathbf{W}^{\text{GeM}}+\alpha\left(\mathbf{A}^{\text{LLM}}-\mathbf{W}^{\text{GeM}}\right)$ (7)
where this formulation automatically patch region weights where the LLM and pre-trained model disagree: if $\mathbf{A}^{\text{LLM}}_{i,j}\approx\mathbf{W}^{\text{GeM}}_{i,j}$, the adjustment term diminishes, thereby avoiding over-emphasis of already salient regions or excessive suppression of already down-weighted areas.
Utilize the $\mathbf{W}^{\text{final}}$ to modulate spatial aggregation while preserving the original mathematical structure of GeM, the attention-weighted GeM pooling is then computed as:
$y_{c}^{\text{att}}=\left(\frac{\sum_{i=1}^{H}\sum_{j=1}^{W}w^{\text{final}}_{i,j}\cdot x_{c,i,j}^{p}}{\sum_{i=1}^{H}\sum_{j=1}^{W}w^{\text{final}}_{i,j}}\right)^{1/p}$ (8)
Spatial weighting in GeM pooling has been demonstrated valuable for image retrieval tasks [46], and our approach extends this concept by leveraging LLM-derived semantic priors at inference time. This enables seamless integration into existing pre-trained VPR models without retraining the backbone networks, while incorporating the powerful reasoning capabilities of LLMs to address challenging recognition scenarios. Given that GeM pooling is widely adopted in state-of-the-art VPR architectures, the proposed VPR-AttLLM integration framework holds substantial potential for enhancing performance across diverse models and challenging visual place recognition tasks.
3.2.3 LLM-Guided Attention for Cluster Aggregation
While GeM pooling performs global spatial aggregation, another parallel paradigm in VPR employs cluster-based aggregation mechanisms, where spatial features are assigned to multiple semantic clusters via soft probabilities, followed by weighted aggregation to construct descriptors. Notable examples include NetVLAD [13] which employ softmax-based soft assignment, as well as its recent variant SALAD [15], which utilizes optimal transport strategy for more robust cluster assignment. Despite architectural differences, these methods share a common aggregation paradigm: for a feature map $\mathbf{X}\in\mathbb{R}^{C\times H\times W}$ with $n$ spatial units, the model computes cluster features $\mathbf{F}\in\mathbb{R}^{l\times n}$ and an assignment matrix $\mathbf{P}\in\mathbb{R}^{m\times n}$ through learned projections, where $l$ denotes the cluster feature dimension and $m$ represents the number of clusters. The standard aggregation computes each cluster’s representation as:
$\mathbf{V}_{:,k}=\sum_{j=1}^{n}P_{k,j}\cdot\mathbf{F}_{:,j},\quad k=1,\ldots,m$ (9)
where $\mathbf{V}_{:,k}\in\mathbb{R}^{l}$ denotes the aggregated feature vector for cluster $k$, $P_{k,j}$ is the assignment probability of spatial unit $j$ to cluster $k$, and $\mathbf{F}_{:,j}\in\mathbb{R}^{l}$ is the feature vector of spatial unit $j$.
Without modifying pre-trained cluster assignments, our LLM attention map $\mathbf{A}^{\text{LLM}}$ is utilized to modulate feature magnitudes, for controlling spatial unit contributions within their assigned clusters. After reshaping $\mathbf{A}^{\text{LLM}}$ to match the feature map’s spatial structure, denoted as $\mathbf{A}^{\text{LLM}}\in\mathbb{R}^{n}$, with blending ratio $\alpha$, the attention-modulated cluster aggregation then becomes:
$\mathbf{V^{att}}_{:,k}=\sum_{j=1}^{n}P_{k,j}\cdot\left(\alpha\cdot A^{\text{LLM}}_{j}\cdot\mathbf{F}_{:,j}\right),\quad k=1,\ldots,m$ (10)
Here $\mathbf{A}^{\text{LLM}}$ serves as a magnitude modulator that differentially scales feature contributions: distinctive regions identified through multimodal reasoning (e.g., unique architectural landmarks) contribute more strongly to their assigned clusters, while generic or ambiguous regions (e.g., common vegetation or occluding objects) exert proportionally reduced influence. This mechanism enables fine-grained discrimination within the same semantic cluster—for instance, emphasizing a unique historical building over a generic modern structure, even though both belong to the "building" cluster. The resulting attention-modulated cluster features are then processed through the original VPR model’s descriptor construction pipeline to obtain the final LLM-guided descriptor.
@@ 4.1.1 SF-XL Dataset and Query Construction
Among existing VPR benchmarks (e.g., Pitts250k [20], Nordland, SF-XL [14]), only SF-XL provides urban-scale ground-truth coverage suitable for city-wide evaluation. The SF-XL reference database contains 2.81 million geo-referenced images sampled from 232k unique locations across San Francisco, making it an ideal testbed for assessing VPR performance at metropolitan scale.
@@
To ensure realism while maintaining geometric consistency essential for VPR evaluation, we employ a carefully designed prompting protocol (detailed in Supplementary Material C) using Gemini-2.0-Flash. Following best practices in synthetic data generation for SVI vision tasks [47, 48, 20], our prompts explicitly instruct preservation of original viewpoints and spatial layouts while introducing only appearance degradations: partial inundation, raindrop blur, surface reflections, and atmospheric haze (Figure 5). Critically, all generated images undergo manual verification to confirm that viewpoints and permanent structural elements remain unmodified—only extreme weather effects are added. This view-preserving protocol ensures that any performance differences stem purely from appearance variations rather than geometric inconsistencies, enabling rigorous controlled evaluation of condition-invariant VPR methods.
4.1.2 HK-URBAN: A New Dataset with Distinctive Urban Morphology
@@
Database construction follows established VPR protocols [20, 14]. Specifically, Google Street View panoramas are sampled at 15 m intervals along all accessible streets within Hong Kong’s central districts. From each panorama, twelve perspective views are rendered with $60^{\circ}$ horizontal field of view (FOV) and $12^{\circ}$ upward pitch to capture building facades and skyline features critical for urban VPR. Query sets mirror those constructed for SF-XL: 100 real social media flood images (hk_flood), 1,000 Mapillary samples (hk_mapi), and 1,000 corresponding simulated flood scenarios (hk_mapi[F]).
@@ 4.2 Implementation Details
To demonstrate the generalizability of the proposed VPR-AttLLM framework, we conduct extensive experiments across diverse state-of-the-art VPR architectures with different backbone networks and aggregation mechanisms. For CNN-based paradigms employing Generalized Mean (GeM) pooling as the final aggregator, we evaluate CosPlace [14] and EigenPlaces [27] using their publicly released 512-dimensional descriptors with VGG16 and ResNet backbones, respectively. For Vision Transformer (ViT)-based paradigms with cluster-based aggregation, we evaluate SALAD [15] in both its full configuration (8192D patch features ++ 256D class token) and compact variant (512D patch features ++ 32D class token), where the first component represents spatially aggregated patch-level descriptors and the second captures global scene semantics from the ViT class token.
@@ 5.1 Effectiveness, model agnosticity, and interpretability of LLM-Guided attention
Beyond quantitative improvements, VPR-AttLLM addresses the undergeneralization problem in discriminative models [21], where VPR systems fail under extreme weather conditions such as flooding. The model-agnostic architecture enables seamless integration with diverse VPR backbones without retraining. The late-stage attention fusion strategy operates directly on feature aggregation pahse for pre-trained global descriptors, making it compatible with both CNN-based models (CosPlace, EigenPlaces with VGG16 and ResNet50) and transformer-based architectures (Salad with DINOv2). This plug-and-play property facilitates rapid deployment and experimentation across existing VPR systems, as evidenced by consistent improvements across all three tested architectures. Furthermore, the framework enhances interpretability by providing natural language justifications for attention weights (e.g., "distinctive curved bay windows contributing significant architectural uniqueness"), allowing practitioners to audit which scene elements drive retrieval decisions—a critical requirement for safety-critical geospatial applications such as emergency response or autonomous navigation.
Finally, this work establishes a novel methodological bridge between urban perception theory and computer vision by operationalizing semantic place understanding through LLM-guided attention. The framework’s ability to identify architecturally distinctive elements (e.g., unique facades, landmark structures) and suppress visually dominant but contextually uninformative regions (e.g., front large but common residential building) empirically validates urban theory about place uniqueness in a computational setting. By encoding human-like spatial reasoning—traditionally studied through cognitive mapping and environmental psychology—into feature weighting mechanisms, these interpretable attention maps generated by VPR-AttLLM offer potential value beyond VPR performance: they could serve as data-driven tools for validating urban design theories about place identity and wayfinding [10], or for auditing training dataset biases in visual geolocation systems. By making explicit which scene elements drive model decisions, LLM-guided frameworks bridge the gap between black-box deep learning and human-interpretable spatial reasoning—a critical step toward trustworthy AI in geospatial applications.
@@ 5.3 Toward Geography-Aware Multimodal VPR Agents
Beyond visual features, multimodal information present in street-level imagery—particularly text elements such as street signs, storefront names, and directional signage—offers complementary localization cues that remain underutilized in current VPR systems. Although such text regions receive high attention weights in LLM-generated attention maps (Figure 7), they cannot be effectively exploited without corresponding geospatial-aware reference databases. Early works in OCR-augmented VPR [49] and recent adavances in LLM-based multimodal reasoning for image geolocalization [36] suggest that fusing visual global descriptors with text embeddings may yield substantial improvements, particularly in urban environments where signage is prevalent and culturally informative.

### v2

@@ 3.2 Attention Integration into VPR models for Inference
We integrate the LLM-generated attention maps into pre-trained VPR models using a modulation strategy that preserves the backbone’s learned structure while injecting semantic guidance. The overall VPR-AttLLM framework is illustrated in Figure 3. Critically, this integration is asymmetric: it applies only to query images at inference time. This allows the system to adapt to challenging scenes (e.g., floods, occlusions) without the prohibitive computational cost of re-processing the entire reference database.
@@ 3.2.1 General Framework Formulation
Let $\mathcal{G}$ be the aggregation module of a pre-trained VPR model that transforms a spatial feature map $\mathbf{X}$ into a global descriptor vector $\mathbf{v}$. We introduce the LLM-generated attention map $\mathbf{A}^{\text{LLM}}$ into this aggregation process, controlled by a blending coefficient $\alpha\in[0,1]$. The attention-modulated inference is defined as:
@@
Here, $\alpha$ acts as a hyperparameter governing the influence of semantic reasoning. When $\alpha=0$, the standard pre-trained behavior is preserved; as $\alpha$ approaches 1, the descriptor formation is increasingly dominated by the LLM’s semantic priors. This formulation allows us to emphasize discriminative regions (e.g., permanent landmarks) and suppress transient noise by scaling the contribution of specific spatial locations in the feature map.
3.2.2 Modulation for GeM Pooling and Cluster Aggregation
We adapt this general principle to two dominant aggregation paradigms: Generalized Mean (GeM) pooling and Cluster-based aggregation.
GeM Pooling Integration. Standard GeM pooling aggregates features by computing a spatially weighted average based on activation intensity. We modify this by blending the model’s native spatial weights $\mathbf{W}^{\text{GeM}}$ with our semantic attention map $\mathbf{A}^{\text{LLM}}$. The final weight map is defined as $\mathbf{W}^{\text{final}}=\mathbf{W}^{\text{GeM}}+\alpha(\mathbf{A}^{\text{LLM}}-\mathbf{W}^{\text{GeM}})$. Consequently, the attention-weighted GeM descriptor is computed as:
$y_{c}^{\text{att}}=\left(\frac{\sum_{i,j}\mathbf{W}^{\text{final}}_{i,j}\cdot x_{c,i,j}^{p}}{\sum_{i,j}\mathbf{W}^{\text{final}}_{i,j}}\right)^{1/p}$ (2)
where $x_{c,i,j}$ represents the feature activation and $p$ is the pooling parameter. This ensures that regions deemed semantically important by the LLM contribute more significantly, correcting cases where the visual model focuses on salient but irrelevant or less identifiable features within given urban context.
Cluster-based Integration. For architectures like NetVLAD or SALAD, which assign local features to semantic clusters before aggregation, we utilize the attention map to modulate feature magnitude. Rather than altering the cluster assignment probabilities—which represent the visual "type" of the feature—we scale the feature vector $\mathbf{F}_{:,j}$ by its corresponding attention value $A^{\text{LLM}}_{j}$ prior to aggregation:
$\mathbf{V^{att}}_{:,k}=\sum_{j=1}^{n}P_{k,j}\cdot\left(\alpha\cdot A^{\text{LLM}}_{j}\cdot\mathbf{F}_{:,j}\right)$ (3)
where $P_{k,j}$ is the assignment probability of spatial unit $j$ to cluster $k$. This acts as a magnitude modulator: a unique building (high attention) will exert a stronger influence on the "building" cluster representation than a generic wall (low attention), enhancing the discriminative power of the resulting global descriptor.
@@ 4.1.1 SF-XL Dataset and Query Construction
Among existing VPR benchmarks (e.g., Pitts250k (torii247Place2018), Nordland, SF-XL (bertonRethinkingVisualGeolocalization2022a)), only SF-XL provides urban-scale ground-truth coverage suitable for city-wide evaluation. The SF-XL reference database contains 2.81 million geo-referenced images sampled from 232k unique locations across San Francisco, making it an ideal testbed for assessing VPR performance at metropolitan scale.
@@
To ensure realism while maintaining geometric consistency essential for VPR evaluation, we employ a carefully designed prompting protocol (detailed in Supplementary Material C. Prompts used for simulating flooding queries) using Gemini-2.0-Flash. Following best practices in synthetic data generation for SVI vision tasks (tremblayRainRenderingEvaluating2021; poravDontWorryWeather2019; torii247Place2018), our prompts explicitly instruct preservation of original viewpoints and spatial layouts while introducing only appearance degradations: partial inundation, raindrop blur, surface reflections, and atmospheric haze (Figure 5). Critically, all generated images undergo manual verification to confirm that viewpoints and permanent structural elements remain unmodified—only extreme weather effects are added. This view-preserving protocol ensures that any performance differences stem purely from appearance variations rather than geometric inconsistencies, enabling rigorous controlled evaluation of condition-invariant VPR methods.
@@
4.1.2 HK-URBAN: A New Dataset with Distinctive Urban Morphology
Database construction follows established VPR protocols (torii247Place2018; bertonRethinkingVisualGeolocalization2022a). Specifically, Google Street View panoramas are sampled at 15 m intervals along all accessible streets within Hong Kong’s central districts. From each panorama, twelve perspective views are rendered with $60^{\circ}$ horizontal field of view (FOV) and $12^{\circ}$ upward pitch to capture building facades and skyline features critical for urban VPR. Query sets mirror those constructed for SF-XL: 100 real social media flood images (hk_flood), 1,000 Mapillary samples (hk_mapi), and 1,000 corresponding simulated flood scenarios (hk_mapi[F]).
@@ 4.2 Implementation Details
To demonstrate the generalizability of the proposed VPR-AttLLM framework, we conduct extensive experiments across diverse state-of-the-art VPR architectures with different backbone networks and aggregation mechanisms. For CNN-based paradigms employing Generalized Mean (GeM) pooling as the final aggregator, we evaluate CosPlace (bertonRethinkingVisualGeolocalization2022a) and EigenPlaces (bertonEigenPlacesTrainingViewpoint2023) using their publicly released 512-dimensional descriptors with VGG16 and ResNet backbones, respectively. For Vision Transformer (ViT)-based paradigms with cluster-based aggregation, we evaluate SALAD (izquierdoOptimalTransportAggregation2024) in both its full configuration (8192D patch features ++ 256D class token) and compact variant (512D patch features ++ 32D class token), where the first component represents spatially aggregated patch-level descriptors and the second captures global scene semantics from the ViT class token.
@@ 5.1 Effectiveness, model agnosticity, and interpretability of LLM-Guided attention
Beyond quantitative improvements, VPR-AttLLM addresses the undergeneralization problem in discriminative models (ilievskiAligningGeneralizationHumans2025), where VPR systems fail under extreme weather conditions such as flooding. The model-agnostic architecture enables seamless integration with diverse VPR backbones without retraining. The late-stage attention fusion strategy operates directly on feature aggregation pahse for pre-trained global descriptors, making it compatible with both CNN-based models (CosPlace, EigenPlaces with VGG16 and ResNet50) and transformer-based architectures (Salad with DINOv2). This plug-and-play property facilitates rapid deployment and experimentation across existing VPR systems, as evidenced by consistent improvements across all three tested architectures. Furthermore, the framework enhances interpretability by providing natural language justifications for attention weights (e.g., "distinctive curved bay windows contributing significant architectural uniqueness"), allowing practitioners to audit which scene elements drive retrieval decisions—a critical requirement for safety-critical geospatial applications such as emergency response or autonomous navigation.
Finally, this work establishes a novel methodological bridge between urban perception theory and computer vision by operationalizing semantic place understanding through LLM-guided attention. The framework’s ability to identify architecturally distinctive elements (e.g., unique facades, landmark structures) and suppress visually dominant but contextually uninformative regions (e.g., front large but common residential building) empirically validates urban theory about place uniqueness in a computational setting. By encoding human-like spatial reasoning—traditionally studied through cognitive mapping and environmental psychology—into feature weighting mechanisms, these interpretable attention maps generated by VPR-AttLLM offer potential value beyond VPR performance: they could serve as data-driven tools for validating urban design theories about place identity and wayfinding (lynchImageCity2008), or for auditing training dataset biases in visual geolocation systems. By making explicit which scene elements drive model decisions, LLM-guided frameworks bridge the gap between black-box deep learning and human-interpretable spatial reasoning—a critical step toward trustworthy AI in geospatial applications.
@@ 5.3 Toward Geography-Aware Multimodal VPR Agents
Beyond visual features, multimodal information present in street-level imagery—particularly text elements such as street signs, storefront names, and directional signage—offers complementary localization cues that remain underutilized in current VPR systems. Although such text regions receive high attention weights in LLM-generated attention maps (Figure 7), they cannot be effectively exploited without corresponding geospatial-aware reference databases. Early works in OCR-augmented VPR (hazelhoffExploitingStreetlevelPanoramic2014) and recent adavances in LLM-based multimodal reasoning for image geolocalization (yinLLMenhancedDisasterGeolocalization2025) suggest that fusing visual global descriptors with text embeddings may yield substantial improvements, particularly in urban environments where signage is prevalent and culturally informative.
@@ Acknowledgements
Data and Code Availability
The data and code of this study are available on request.